OpenStudy (anonymous):

state the domain and range of each of the following functions f(x)=1/x+3

OpenStudy (anonymous):

B)g(x)=sqrt x+6 C)h(x)=x^3+2x+5

OpenStudy (jhannybean):

The domain is found when the function \(x+3 = 0\)

OpenStudy (jhannybean):

Typo.

OpenStudy (anonymous):